OpenStudy (anonymous):

a right triangles hypotenuse is 6m and one legs is 2 square root 5 m. find the length of the other leg

OpenStudy (cwrw238):

use pythagoras 6^2 = (2sqrt5)^2 + x^2 x = sqrt(36 - 20)

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow (2\sqrt{5})^2 + x^2 = 6^2 }\)

Parth (parthkohli):

This is the Pythagorean theorem

OpenStudy (anonymous):

\[b=\sqrt{c^2-a^2}\] where c is the hypotenuse and a is the leg as parth mention pythagorean theorem \[c^2=a^2+b^2\] where c=hypotenuse a, b are the legs